OpenStudy (anonymous):

1. The combustion of a sample of butane, C4H10 (lighter fluid), produced 2.46 grams of water. 2 C4H10 + 13 O2 -------> 8 CO2 + 10 H2O (a) How many moles of water formed? (b) How many moles of butane burned? (c) How many grams of butane burned? (d) How much oxygen was used up in moles? In grams?

OpenStudy (abb0t):

Using what you're given, meaning \(\sf \color{red}{2.46~g~H_2O}\) multiply that value by the \(molar~mass\) of H\(_2\)O. Once, you've done that, you can get moles of water formed.

OpenStudy (abb0t):

Essentially, what you're doing here is canceling out units, this is called dimensional analysis. very similar to what you learned in algebra: \(\sf \color{}{\frac{g}{1}\times \frac{mole}{g}}\), and notice how the units for \(g\) cancel out? You're left with moles of water.